OpenStudy (anonymous):

Help please! :) Write x^2 - 8x + 13 = 0 in the form (x - a)^2 = b, where a and b are integers. (x - 4)^2 = 3 (x - 3)^2 = 2 (x - 2)^2 = 1 (x - 1)^2 = 4

OpenStudy (dumbcow):

sorry -take half of "b" coefficient, then square it - subtract from "c" coefficient - factor perfect square -leaving --> (x- b/2)^2 + (c - b^2/4)

OpenStudy (anonymous):

the answer is first option (x-4)^2 =3 we need to get 8x through the squaring proceedure!! which can be obtained only by squaring (x-4) = X^2 ((-2*4*x)) +16 =3 the three taken to LHS becomes -3 and 16-3 will give 13

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

if it is an objective question (only answer matters) you can get away with the short cut!! if it is a subjective one then then youll need to re trace your steps from the answer!! Good luck!!